OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

Vocaloid (vocaloid):

speed = distance/time car A speed = x/2.4 car B speed = x/4 x/24 = (x/4) + 22 solve for x

OpenStudy (anonymous):

ok that kinda helps but i still don't get it...

Vocaloid (vocaloid):

we want to find the speed of car A, right? if we let x = distance, then speed of car A = x/2.4, with me so far?

OpenStudy (anonymous):

yes i'm with you

Vocaloid (vocaloid):

and speed of car B = x/4, right?

OpenStudy (anonymous):

right

Vocaloid (vocaloid):

the problem tells us that car A travels 22 mph faster than car B, so car A speed = car B speed + 22, right?

OpenStudy (anonymous):

i think so

Vocaloid (vocaloid):

so we substitute car A speed = x/2.4 and car B speed = x/4 to get x/2.4 = (x/4) + 22 and solve for x

OpenStudy (anonymous):

ok

Vocaloid (vocaloid):

so, what does x = ?

OpenStudy (anonymous):

is it 132 ?

Vocaloid (vocaloid):

yes! so x = distance = 132 miles all we need to do now is find speed of car A = x/2.4 = 132/2.4 = 55mph

OpenStudy (anonymous):

yay thanks for the help!